Bayside insurance offers two health plans. Under plan A, Giselle would pay the first $50 of her medical bills, plus 25% of the rest. Under plan B, Giselle would pay the first $70, but only 20% of the rest. For whatamount of medical bills will plan B save Giselle money? Assume she has over $70 in bills.

Answer by mananth(16946)

Bayside insurance offers two health plans. Under plan A, Giselle would pay the first $50 of her medical bills, plus 25% of the rest. Under plan B, Giselle would pay the first $70, but only 20% of the rest. For whatamount of medical bills will plan B save Giselle money? Assume she has over $70 in bills.
PLAN A
Let bill be x
payment = x-50+0.25(x-50)
payment = x-50+0.25x-12.50
payment = 0.75x-62.50
...
PLAN B
payment = x-70 +0.2(x-70)
payment = x-70 +0.2x-14
payment = 0.8x-84
PLAN B < PLAN A
0.8x-84<0.75x-62.50
0.8x-0.75x<84-62.50
0.05x<21.50
/0.05
x<21.50/0.05
x<430
for Plan B to save money bill value must be <$430